OpenStudy (anonymous):

Find the "LCM" of z^2-4 and z^2+5z+6.

jimthompson5910 (jim_thompson5910):

z^2-4 = (z-2)(z+2) z^2+5z+6 = (z+2)(z+3) Now collect the unique factors and multiply them out to get the LCM of (z-2)(z+2)(z+3) Optionally, you can expand and simplify, but it's not necessary.

OpenStudy (anonymous):

Cool! Thank you! :)